Please help me with this question:
Let f(x)=3-2x and g(x)=sqrt(x) and h(x)=3x^2+2
Find (f+h)(5)
Thank you very much! :)

Answer by tommyt3rd(5050)

(f+h)(5)=
f(5)+h(5)=
[3-2(5)]+[3*(5)^2+2]=
70